1. The term structure of interest rates refers to the relationship between _____.

a bond's time to maturity and its coupon rate

a bond's age since issue and its coupon rate

a bond's age since issue and its yield

a bond's time to maturity and its yield.

2. The yield on 12-month treasury bills is 1.4% and the yield on 2-year treasury STRIPS is 2%.

a. What is the implied 1-year forward rate one year from now?

3. The term structure shows the following Treasury spot rates:

Treasury spot interest rates are as follows:

Maturity (years) 1 2 3 4
Spot rate (EAR) 1.9% 2.8% 3.6% 4.5%

What is the price of a risk-free zero-coupon bond with 3 years to maturity and a face value of $1,000 (in $)?
Answer #1

1.
a bond's time to maturity and its yield.

2.
a. What is the implied 1-year forward rate one year from now?
=1.02^2/1.014-1=2.60%

3.
What is the price of a risk-free zero-coupon bond with 3 years to maturity and a face value of $1,000 (in $)?
=1000/(1+3.6%)^4
=868.0824543